# Supercoiling DNA with a free end

### v1

@@ 4 The three-dimensional polymer case
Fig. 6(a) shows the same conformations at different times of Fig. 1(b), at $\tau_{a}=6$, now with beads colored in blue in the plectonemic regions, and red otherwise, using the aforementioned detection criteria based on the writhe threshold. Fig. 6(b) shows the contact map, computed by quantifying the spatial proximity of all pairs of monomers along the polymer chain. A contact between monomers $i$ and $j$ is defined whenever their mutual distance $d_{ij}=\sqrt{(\textbf{r}_{i}-\textbf{r}_{j})^{2}}$ falls below a chosen cutoff value $d_{cut}=6\sigma$. The resulting two-dimensional matrix provides a global view of the three-dimensional organization of the polymer: diagonal elements represent local backbone connectivity. At the same time, off-diagonal clusters of contacts reflect plectonemic structures. Fig.6(c) shows the local writhe as a function of the chain index, respectively.
@@
Overall, Fig. 6 highlights that a first plectoneme starts forming in the middle of the chain, as indicated by a larger peak in the local writhe and a larger off-diagonal band in the contact maps, and then grows over time until it eventually spans the entire chain. At intermediate times, we also observe the formation of additional smaller plectonemes, which merge with the larger one dynamically, similarly to what was observed in [16]. This picture becomes clearer when examining the kymograph in Fig. 7, which plots regions of the chain with plectonemes, coloring the largest one in red and the smaller ones in cyan. The first plectoneme starts forming and growing at around $5l_{b}$ from the point of active torque application, while the others form immediately in the vicinity but do not grow and are eventually merged with the larger one.
A striking feature is that, once formed, the torque effectively pins the plectoneme close to the point of active torque application, and for the smallest chain, effectively spans the entire chain. Moreover, a positive $\tau_{a}$ generates a writhe which is negative in value.
@@
Figure 8: The first and second rows show results from the statistical field theory, for the plectonemic and swollen phases, respectively. (a) Total twist and writhe, $\rm Tw$} and $\rm Wr$}, as a function of time. The dashed line indicates $t^{0.5}$. (b) Local twist density, $\rho_{\rm tw}$, and centerline writhe density, $\rho_{\rm wr}$, as a function of the chain contour length $s$. Parameters for the field theory model are: $\alpha=0.1$, $\beta=0.2$, $\gamma=0.15$, $\kappa=0.1$, and $\sigma_{0}=1$. For the swollen phase, we set $D_{\rm Tw}=100$, $D_{\rm Wr}=1$, and ${\rm Tw}_{0}=0.2$. For the plectonemic phase we set $D_{\rm Tw}=80$, $D_{\rm Wr}=20$, and ${\rm Tw}_{0}=1.2$.
Field theory for supercoiling transport. The qualitative trends observed up to now can be rationalized by using a field theory for supercoiling transport, which is based on the continuum equations of motion proposed in [16] for the coupled reaction-diffusion dynamics of local twist density, $\rho_{\rm tw}(s,t)$, and centerline writhe density, $\rho_{\rm wr}(s,t)$, with $s$ the chain contour length, generalized to the boundary conditions which are appropriate for the setup we consider. These equations read as follows,
@@
$g(\rho_{\rm tw},\rho_{\rm wr})$ $=$ $\alpha\rho_{\rm tw}-\beta\rho_{\rm wr}+\gamma\Theta(|\rho_{\rm tw}+\rho_{\rm wr}|-\sigma_{0})$
In Eqs. (21), $D_{\rm tw}$ and $D_{\rm wr}$ denote respectively the twist and the writhe effective diffusion coefficients. At the same time, $\kappa$ is a constant which contributes to the surface tension of a plectoneme and the width of the interface between a plectoneme and a straight segment. Note that the equation of $\rho_{\rm tw}$ is of the same form as the one we found as Eq.15, with the addition of the function $g(\rho_{\rm tw},\rho_{\rm wr})$, a reaction-like term which controls the mutual interconversion between local twist and local writhe. The active torque term in this case is imposed as a boundary condition. In the form we have chosen for $g$ [16], $\alpha$ and $\beta$ determine the rates of conversion from writhe to twist and vice versa, whereas $\gamma$ modulates this interconversion when the local supercoiling $\sigma$ exceeds $\sigma_{0}$: if $\alpha\geq\beta\geq\gamma$, a chain with uniform twist is locally unstable to plectoneme formation as soon as $\sigma>\sigma_{0}$. With respect to the field theory proposed in [16], we disregard noise here – this is reasonable, as we see little evidence of plectoneme diffusion along the chain in the simulations. Importantly, instead of periodic boundary conditions considered in [16], here we consider Dirichlet-like boundary conditions for the twist and writhe fields, with $\rho_{\rm tw}(s=0)={\rho_{\rm tw}}_{0}$ at the end which is subject to the active torque, and $\rho_{\rm tw}(s=N\sigma)=0$ at the open end: these are equivalent to including a $\tau\delta^{\prime}(x)$ forcing to the twist equation. For the writhe field, simulations suggest that $\rho_{\rm wr}$ should be 0 at both ends.
@@
Fig. 8(a) shows the total accumulated twist, $\rm Tw$} and writhe, $\rm Wr$} for the plectonemic and swollen cases, as a function of time, obtained integrating $\rho_{\rm tw}$ and $\rho_{\rm wr}$ over the entire chain length. First, in both cases, the twist accumulates over time approximately diffusively, as $t^{0.5}$, while the effective exponent of writhe accumulation varies over time. Twist and writhe reach a steady state eventually, with the writhe becoming smaller than the twist in the swollen phase.
Fig. 8(b) shows instead the steady-state profiles of $\rho_{\rm tw}$ and $\rho_{\rm wr}$ as a function of the contour length $s$. In the swollen phase, the local twist profile decays linearly, compatible with the expectation of Eq.12. In the plectonemic phase, we observe instead a non-linear twist profile reminiscent of that found in molecular dynamic simulations, which corresponds to a slowdown in supercoiling transport due to plectoneme formation. The profile of $\rho_{\rm wr}$ shows again a linear decrease for $D_{\rm Tw}\gg D_{\rm Wr}$, while for $D_{\rm Tw}\sim D_{\rm Wr}$ a bump emerges, which should correspond to a region where the plectoneme establishes. Again, this region is adjacent to the point of active torque application. Note that in the plectonemic phase, the mean field theory predicts also a small non-monotonic region in the twist profile close to the plectoneme boundary, which we do not found in simulations, possibly due to the neglect of noise in the field theory which disallows fluctuation in the position of such boundary, or due to the limited size of the chains considered in molecular dynamics simulations, as plectonemes span the entire chain.
@@ Conclusions
At sub-threshold torques ($\tau_{a}<\tau_{th}$, with $\tau_{th}\sim 2$), the chain is swollen and follows a behavior which is similar to that of a chain constrained to be linear: the intrinsic angular velocity is about constant along the chain and scale approximately as $\omega_{L}\sim\tau_{a}/N$, the local twist decays linearly along the chain, starting from the end where $\tau_{a}$ is applied, and the gyration radius does not change with respect to the value at $\tau_{a}=0$. Dynamically, the accumulation of total twist follows a diffusive law $\langle{\rm Tw}\rangle\sim t^{0.5}$ until saturating at the chain’s torsional capacity, while the total bending remains mostly constant.
@@
Taken together, these results demonstrate that there is a non-equilibrium torque-driven transition from a swollen phase with uniform torsional diffusion to localized supercoiling in the plectonemic phase, accompanied by an associated non-linear twist profile in the steady state. This transition is fundamentally distinct from the buckling of a tethered DNA under twist, which typically requires the presence of a stretching force [12, 27]. The free-chain-end geometry that we have considered could be recreated in the lab through single-molecule experiments, for instance, by studying transcription on DNA templates in solution, if the setup of [23] could be extended in the future to work on untethered DNA molecules. Interestingly, the results in [23] show that an RNA polymerase may be viewed as a torque-generated machine even if its rotational motion is not hindered. Because polymerases can exert a torque of a few $k_{B}T$ [22, 37], it is possible that the torque-induced supercoiling with a free end we have studied is relevant in a biological context, where DNA or chromatin is transcribed [3]. Note that, of course, a transcribing polymerase would exert a torque dipole on the polymer [3], resulting in the formation of twin domains. Therefore, our analysis in practice applies to the behaviour of one of the two supercoiling domains that would form.

### v2

@@ 4 The three-dimensional polymer case
Fig. 6(a) shows the same conformations at different times of Fig. 1(b), at $\tau_{a}=6$, now with beads colored in blue in the plectonemic regions, and red otherwise, using the aforementioned detection criteria based on the writhe threshold. Fig. 6(b) shows the contact map, computed by quantifying the spatial proximity of all pairs of monomers along the polymer chain. A contact between monomers $i$ and $j$ is defined whenever their mutual distance $d_{ij}=\sqrt{(\textbf{r}_{i}-\textbf{r}_{j})^{2}}$ falls below a chosen cutoff value $d_{cut}=6\sigma$. The resulting two-dimensional matrix provides a global view of the three-dimensional organization of the polymer: diagonal elements represent local backbone connectivity. At the same time, off-diagonal clusters of contacts reflect plectonemic structures. Fig.6(c) shows the local writhe as a function of the chain index, respectively, where we identify plectoneme regions here as those where ${\rm Wr_{i}}\leq-0.5$. Note that a positive $\tau_{a}$ generates a writhe which is negative in value.
@@
Overall, Fig. 6 highlights that a first plectoneme starts forming near the end where torque is applied, see peak in the local writhe around bead 80 and a larger off-diagonal band in the contact maps; then the plectoneme grows over time until it eventually spans the entire chain. At intermediate times, we also observe the formation of additional smaller plectonemes, which merge with the larger one dynamically, similarly to what was observed in [16]. This picture becomes clearer when examining the kymograph in Fig. 7, which plots regions of the chain with plectonemes, coloring the largest one in red and the smaller ones in cyan. The first plectoneme starts forming and growing at around $4l_{b}$ from the point of active torque application, while the others form immediately in the vicinity but do not grow and are eventually merged with the larger one.
Note that in our setup, at any given time, the sum of the total twist and writhe is not constrained to be constant. In fact, Fig. 8 shows Tw and Wr, averaged over different realizations, for $\tau_{a}=1,6$ and a 22kbp long chain. In both cases, Tw reaches a steady-state value faster than Wr. Notably, at $\tau_{a}=1$, in the swollen phase, Tw is much larger than Wr, while in the plectonemic phase at $\tau_{a}=6$ they reach a similar steady-state value.
We can also characterize the local writhe profile, ${\rm Wr}_{i}$. Fig. 9 shows these profiles, for $\tau_{a}=6$, averaged over different conformations at steady state. We observe that the curves are asymmetric. Increasing the chain length, we observe two phenomena: the peak position moves towards the chain end where torque is applied, while at the opposite end we have a small portion of the chain free from plectonemes. Thus, a striking feature is that, once formed, the torque effectively pins the plectoneme close to the point of active torque application.
@@
Figure 8: Total twist and writhe, Tw and Wr (Eqs. 18 and 21), as a function of time, for a 22kbp long chain and $\tau_{a}=6$, averaged over several realizations. The values displayed in panels (a) and (b) are for $\tau_{a}=6$ and $1$, corresponding to the plectonemic and swollen phases, respectively.
Figure 9: Local writhe $\rm Wr_{i}$} as a function of the chain fraction $i/N$, averaged over different conformations at steady state, shown for different chain lengths.
Figure 10: The first and second rows show results from the statistical field theory, for the plectonemic and swollen phases, respectively. (a) Total twist and writhe, $\rm Tw$} and $\rm Wr$}, as a function of time. The dashed line indicates $t^{0.5}$. (b) Local twist density, $\rho_{\rm tw}$, and centerline writhe density, $\rho_{\rm wr}$, as a function of the chain contour length $s$. Parameters for the field theory model are: $\alpha=0.1$, $\beta=0.2$, $\gamma=0.15$, $\kappa=0.1$, and $\sigma_{0}=1$. For the swollen phase, we set $D_{\rm tw}=100$, $D_{\rm wr}=1$, and $\rho_{{\rm tw}_{0}}=0.2$. For the plectonemic phase we set $D_{\rm tw}=80$, $D_{\rm wr}=20$, and $\rho_{{\rm tw}_{0}}=1.2$.
Field theory for supercoiling transport. The qualitative trends observed up to now can be rationalized by using a field theory for supercoiling transport, which is based on the continuum equations of motion proposed in [16] for the coupled reaction-diffusion dynamics of local twist density, $\rho_{\rm tw}(s,t)$, and centerline writhe density, $\rho_{\rm wr}(s,t)$, with $s$ the chain contour length, generalized to the boundary conditions which are appropriate for the setup we consider. These equations read as follows:
@@
$g(\rho_{\rm tw},\rho_{\rm wr})$ $=$ $\alpha\rho_{\rm tw}-\beta\rho_{\rm wr}+\gamma\Theta(|\rho_{\rm tw}+\rho_{\rm wr}|-\sigma_{0}).$
In Eqs. (22), $D_{\rm tw}$ and $D_{\rm wr}$ denote respectively the twist and the writhe effective diffusion coefficients. At the same time, $\kappa$ is a constant which contributes to the surface tension of a plectoneme and the width of the interface between a plectoneme and a straight segment. Note that the equation of $\rho_{\rm tw}$ is of the same form as the one we found as Eq.15, with the addition of the function $g(\rho_{\rm tw},\rho_{\rm wr})$, a reaction-like term which controls the mutual interconversion between local twist and local writhe. The active torque term in this case is imposed as a boundary condition. In the form we have chosen for $g$ [16], $\alpha$ and $\beta$ determine the rates of conversion from writhe to twist and vice versa, whereas $\gamma$ modulates this interconversion when the local supercoiling $\sigma_{sc}$ exceeds $\sigma_{0}$: if $\alpha\geq\beta\geq\gamma$, a chain with uniform twist is locally unstable to plectoneme formation as soon as $\sigma_{sc}>\sigma_{0}$. With respect to the field theory proposed in [16], we disregard noise here – this is reasonable, as we see little evidence of plectoneme diffusion along the chain in the simulations. Importantly, instead of periodic boundary conditions considered in [16], here we consider Dirichlet-like boundary conditions for the twist and writhe fields, with $\rho_{\rm tw}(s=0)={\rho_{\rm tw}}_{0}$ at the end which is subject to the active torque, and $\rho_{\rm tw}(s=N\sigma)=0$ at the open end: these are equivalent to including a $\tau\delta^{\prime}(x)$ forcing to the twist equation. For the writhe field, simulations suggest that $\rho_{\rm wr}$ should be 0 at both ends. Note also that starting from Eq. 22 one can obtain an effective free energy mathematically equivalent to the two-state free energy introduced by Marko [14], see Ref. [16].
@@
Fig. 10(a) shows the total accumulated twist, $\rm Tw$} and writhe, $\rm Wr$} for the plectonemic and swollen cases, as a function of time, obtained integrating $\rho_{\rm tw}$ and $\rho_{\rm wr}$ over the entire chain length. First, in both cases, the twist accumulates over time approximately diffusively, as $t^{0.5}$, while the effective exponent of writhe accumulation varies over time. Twist and writhe reach a steady state eventually; the swollen and plectonemic phases are distinguishible in this model by the fact that in the first case the writhe is smaller than the twist, while in the second case they become of the same order. This results parallel the one of Fig. 8 for the polymer model.
Fig. 10(b) shows instead the steady-state profiles of $\rho_{\rm tw}$ and $\rho_{\rm wr}$ as a function of the contour length $s$. In the swollen phase, the local twist and writhe profiles decay linearly, with the former being more dominant and compatible with the expectation of Eq.12. In the plectonemic phase, we observe instead a non-linear twist profile reminiscent of that found in molecular dynamic simulations, which corresponds to a slowdown in supercoiling transport due to plectoneme formation. The profile of $\rho_{\rm wr}$ shows again a linear decrease for $D_{\rm tw}\gg D_{\rm wr}$, while for $D_{\rm tw}\sim D_{\rm wr}$ a bump emerges, which should correspond to a region where the plectoneme establishes. Again, this region is adjacent to the point of active torque application, and parallels the one of Fig. 9, highlighting that the plectonemic phase is characterized by a writhe density contribution dominant with respect to the twist density. Note that in the plectonemic phase, the mean field theory predicts also a small non-monotonic region in the twist profile close to the plectoneme boundary, which we do not found in simulations, possibly due to the neglect of noise in the field theory which disallows fluctuation in the position of such boundary, or due to the limited size of the chains considered in molecular dynamics simulations, as plectonemes span the entire chain.
@@ Conclusions
At sub-threshold torques ($\tau_{a}<\tau_{th}$, with $\tau_{th}\sim 2$), the chain is swollen and follows a behavior which is similar to that of a chain constrained to be straight: the intrinsic angular velocity is about constant along the chain and scale approximately as $\omega_{L}\sim\tau_{a}/N$, the local twist decays linearly along the chain, starting from the end where $\tau_{a}$ is applied, and the gyration radius does not change with respect to the value at $\tau_{a}=0$. Dynamically, the accumulation of total twist follows a diffusive law $\langle{\rm Tw}\rangle\sim t^{0.5}$ until saturating at the chain’s torsional capacity, while the total bending remains mostly constant.
@@
Taken together, these results demonstrate that there is a non-equilibrium torque-driven transition from a swollen phase with uniform torsional diffusion to localized supercoiling in the plectonemic phase, accompanied by an associated non-linear twist profile in the steady state. This transition is fundamentally distinct from the buckling of a tethered DNA under twist, which typically requires the presence of a stretching force [31, 12]. The free-chain-end geometry that we have considered could be recreated in the lab through single-molecule experiments, for instance, by studying transcription on DNA templates in solution [22, 23], via optical or magnetic tweezers [24, 25], or even via electrosmotic flow [26, 27]. Interestingly, the results in [23] show that an RNA polymerase may be viewed as a torque-generated machine even if its rotational motion is not hindered. Because polymerases can exert a torque of a few $k_{B}T$ [22, 42], it is possible that the torque-induced supercoiling with a free end we have studied is relevant in a biological context, where DNA or chromatin is transcribed [3]. Note that, of course, a transcribing polymerase would exert a torque dipole on the polymer [3], resulting in the formation of twin domains. Therefore, our analysis in practice applies to the behaviour of one of the two supercoiling domains that would form.